When santy and lorna were married, his age was 3/2 of her age. If on their golden wedding anniversary, Santy's age will be 8/7 of lorna's age, how old will each of them be on their golden anniversary. Show algebraic solution
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

when they got married, he was 30 and she was 20.
the ratio of their ages was 3/2.

50 years later, he was 80 and she was 70.
the ratio of their ages was 8/7.

to solve this algebraically, you can do the following.

let a = his age when they got married.
let b = her age when they got married.

you get a/b = 3/2.

solve for a to get a = 3/2 * b.

50 years later his age is a + 50.
50 years later her age is b + 50.

you get (a + 50) / (b + 50) = 8/7

solve for (a + 50) to get (a + 50) = 8/7 * (b + 50).

solve for a to get a = 8/7 * (b + 50) - 50

this means that 3/2 * b = 8/7 * (b + 50) - 50 because they are both equal to a.

multiply both sides of this equation by 2/3 and you get:

b = 2/3 * 8/7 * (b + 50) - 2/3 * 50

since 2/3 * 8/7 is equal to 16/21 and since 2/3 is equal to 14/21, this equation becomes:

b = 16/21 * (b + 50) - 14/21 * 50

simplify this to get:

b = 16/21 * b + 16/21 * 50 - 14/21 * 50

combine like terms to get:

b = 16/21 * b + 2/21 * 50

subtract 16/21 * b from both sides of this equation to get:

b - 16/21 * b = 2/21 * 50

since b = 21/21 * b, this becomes:

21/21 * b - 16/21 * b = 2/21 * 50

combine like terms to get 5/21 * b = 2/21 * 50

multiply both sides of this equation by 21 to get 5 * b = 2 * 50

simplify to get 5 * b = 100

divide both sides of this equation by 5 to get b = 20

since a = 3/2 * b, you get a = 30

50 years later you get a + 50 = 80 and b + 50 = 70

the ratio of their ages is 30/20 = 3/2 when they got married.

the ratio of their ages is 80/70 = 8/7 on their golden wedding anniversary.





Answer by ikleyn(52776)

.
Solved in
https://www.algebra.com/algebra/homework/word/age/Age_Word_Problems.faq.question.1062746.html

https://www.algebra.com/algebra/homework/word/age/Age_Word_Problems.faq.question.1062746.html